OpenStudy (anonymous):

The area of a rectangular garden is 54 square feet. If the width of the garden is 1.2 feet, what is its length, in feet? A. 4.5 B. 16 C. 45 D. 64.8

OpenStudy (nurali):

A=wl A=54 ft^2 and w=1.2 ft l=A/w=54/1.2=45ft

OpenStudy (anonymous):

thanks